# The FAST-SETI Milky Way Globular Cluster Survey I: A Pilot Multibeam On-the-Fly Search of Five Globular Clusters at L-Band

### v1

@@ 1 Introduction
The search for technosignatures—observable indicators of technology developed by extraterrestrial intelligence—has long emphasized radio transmissions, particularly ultra–narrowband features (Hz–level) that are widely treated as hallmarks of engineered emitters rather than natural astrophysical processes. This logic underpins many modern radio SETI programs, which prioritize searches for narrowband carriers across GHz bands (e.g., Siemion2013Kepler; Enriquez2017BL692; Price2020BL1327).
@@
Despite decades of effort, no confirmed extraterrestrial signal has been identified. A central challenge is the ubiquity of terrestrial radio–frequency interference (RFI), which can closely mimic sky–localized technosignatures and inflate candidate lists by orders of magnitude before vetting (e.g., Enriquez2017BL692; Price2020BL1327). These experiences motivate observing strategies and pipelines that discriminate RFI in situ and at scale.
@@
To that end, a MultiBeam Point-Source Scanning (MBPS) strategy has been proposed and tested with the Five-hundred-meter Aperture Spherical radio Telescope (FAST). MBPS exploits a multibeam receiver and controlled slews so that a sky-localized signal appears only in the beam presently aligned with the source, whereas broad RFI typically enters multiple beams simultaneously. In a FAST trial, MBPS logic rejected all initial narrowband candidates as RFI, illustrating the approach’s power to suppress persistent false positives (Huang2023MBPS).
@@
FAST is exceptionally well suited to such strategies. Its 19-beam L-band receiver (1.05–1.45 GHz) and low system temperature deliver outstanding sensitivity for technosignature searches (Jiang2020FAST; Nan2011; Li2016). The multibeam architecture is integral to MBPS and enables simultaneous on-source/off-source comparisons within a single pointing pattern.
@@
Beyond nearby stars, there is growing interest in extending searches to diverse environments. Globular clusters (GCs) are ancient, gravitationally bound stellar systems with $10^{5}$$10^{6}$ members; their high stellar densities could shorten communication and travel times within a putative technological network, yet crowding and typically low metallicity may inhibit planet formation and long-term orbital stability (FischerValenti2005; BeerKingPringle2004; Sigurdsson2003; DiStefano2016). From an observational perspective, GC demographics and structural parameters are well compiled in the Harris catalog and updates, which inform experiment design (harris1996).
@@
This work presents the first FAST SETI survey dedicated to GCs and the first to employ MBPS as the primary observing strategy. We target NGC 6171 (M107), NGC 6218 (M12), NGC 6254 (M10), NGC 6838 (M71), and IC 1276 (Pal 7). Our pilot survey demonstrates MBPS in extended, dense stellar fields and places sensitive upper limits on powerful radio transmitters within these systems. Section 2 summarizes observations and configuration; Section 3 details search methods, RFI assessment, and sensitivity; Section 5 discusses implications and prospects for future GC technosignature searches.
2 Observations
2.1 Target Selection
@@
Rationale. Globular clusters (GCs) are ancient, dense stellar systems that offer both astrophysical motivation and practical advantages for technosignature searches. Theoretical work has argued that, despite high stellar encounter rates in cluster cores, there exist substantial “sweet-spot” regions in many GCs where habitable-zone orbits can remain dynamically stable over gigayear timescales; the short interstellar separations in such regions would also greatly facilitate interstellar communication and travel for any advanced civilization (DiStefano2016). At the same time, the well-established planet–metallicity correlation suggests that stars drawn from the metal-richer tail of the GC distribution should be more likely to host giant planets (and perhaps complex planetary architectures) than their more metal-poor counterparts (FischerValenti2005). Empirically, planets can exist in GCs: the hierarchical triple in M4 (PSR B1620-26 with a planetary-mass companion) provides a long-discussed proof-of-concept that planetary bodies can survive, or be (re)assembled, in low-metallicity cluster environments (Sigurdsson2003; BeerKingPringle2004). Together, these considerations motivate a focused SETI experiment on relatively metal-rich Milky Way GCs.
@@
    Metallicity (planet formation prior). We prioritized clusters toward the metal-rich end of the Galactic GC distribution, where the prior probability for giant-planet occurrence is higher (FischerValenti2005). In particular, M71 and Pal 7 are among the more metal-rich Milky Way GCs, while M10/M12/M107 are of intermediate metallicity. All metallicities quoted in this paper are drawn uniformly from the Harris catalog (1996; 2010 edition) to ensure internal consistency (harris1996).
@@
    Proximity (EIRP sensitivity). For a flux-limited narrowband search, the minimum detectable equivalent isotropic radiated power scales as $d^{2}$. We therefore favored nearby clusters (heliocentric distances $\sim$4–6.5 kpc), improving our EIRP limits for any continuous narrowband transmitters (harris1996).
@@
    Stellar population size (trial factor). Integrated absolute magnitudes ($M_{V}$) from the Harris catalog were used as proxies for total stellar population. Brighter clusters (e.g., M10, M12, M107 with $M_{V}\sim-7$) host larger numbers of stars, increasing the chance of an n=1n{=}1 technosignature source within the beam at any instant (harris1996).
@@
Scientific expectations. From an SETI perspective, this sample spans a scientifically useful range of cluster environments while keeping the priors favorable. The inclusion of comparatively metal-rich disk/bulge clusters (M71, Pal 7) is directly aligned with planet-occurrence evidence (FischerValenti2005), and the presence of the PSR B1620-26 planetary system in M4 demonstrates that planetary companions (and complex dynamical assembly channels) are realizable in GCs (Sigurdsson2003; BeerKingPringle2004). Theoretical work indicates that, even in dynamically active systems, large volumes outside the densest cores can host long-lived habitable zones (DiStefano2016). Observationally, all five targets are close enough that FAST’s sensitivity yields stringent EIRP limits for continuous, narrowband carriers (see §3.3), and their declinations place them in FAST’s high-gain sky. For reproducibility and uniformity, the equatorial coordinates, distances, metallicities, and $M_{V}$ values used here are taken from the Harris (2010 edition) catalog (harris1996); the specific parameter values and the observation windows are listed in Table 1 and Table 2.
Table 1: Target Globular Clusters and Key Properties (from harris1996)
@@
Note. — Harris (1996; 2010 edition) values. Distances are heliocentric; metallicities in the Zinn–West scale. Angular sizes are from 1989Sci
@@
2.2 FAST Instrument, MBOTF Scanning, and Geometry
@@
and the separation between adjacent on–target beam centers ($\sim$5.74′) implies $\sim$115 s between consecutive crossing midpoints at the commanded scan rate. These deterministic spacings (in angle and in time), together with the two-pass complementary coverage, are the geometric constraints that underpin our later multibeam verification (§3.2).
For subsequent analysis across the full 19-beam footprint, the array tracks four adjacent declination stripes during the raster. We therefore refer to four beam groups (BGs), A–D, each denoting one stripe traced by a subset of beams in a pass. The central stripe (BG–C) contains the five-beam line that passes the cluster core; the adjacent stripes (BG–A/B/D) provide complementary illumination above and below BG–C. This stripe-based view is purely geometric and does not change the instrument configuration; it serves to organize timing/order expectations and coverage accounting in the cross–verification stage (§3.2).
3 Data Analysis
3.1 Primary Search Configuration and Gating
@@
Spectra are analyzed at $\sim$7.5 Hz frequency resolution and $\sim$10 s integration. Linear polarizations (XX and YY) are processed independently over a drift-rate range $|{\dot{\nu}}|\leq 4$ Hz s-1, and events with ${\rm S/N}\geq 10$ are recorded. The resulting per-beam hit lists contain the BG label, pass index, gOTW identifier, boresight offset from the mask, and boundary flags marking hits within a few seconds of gOTW edges. Geometry-based validation is applied afterward (§3.2). A conservative HPBW corresponding to the widest beam in 1.05–1.45 GHz is used to ensure uniform gating across the band.
3.2 MBPS Cross-Verification Using Beam Groups
@@ 3.3 Sensitivity and EIRP
3.4 Choice of Maximum Drift Rate
@@
    Astrophysical plausibility. Earth’s orbital and rotational accelerations yield drifts $\lesssim 0.2~\mathrm{Hz~s^{-1}}$ at L-Band (Siemion2013Kepler). A transmitter on an Earth-like planet would plausibly add a similar contribution, placing realistic drifts in the $\sim 0.3$$0.5~\mathrm{Hz~s^{-1}}$ range. Our ceiling exceeds this by nearly an order of magnitude, comfortably encompassing Earth-like and moderately more extreme kinematics.
@@
    Algorithmic feasibility. With $\Delta\dot{\nu}=0.125~\mathrm{Hz~s^{-1}}$, $\pm 4~\mathrm{Hz~s^{-1}}$ implies $\sim 65$ trials per scan, a computationally efficient load for multi-hour FAST datasets. Extending to the “200 nHz” normalized envelope ($\dot{\nu}_{\max}\!\sim\!280~\mathrm{Hz~s^{-1}}$ at 1.4 GHz; Sheikh2020ETZ) would inflate trials by a factor of $\sim 70$, substantially increasing runtime and false-positive candidates without clear gains for 60 s scans.
@@
    Comparative precedent. Our choice aligns with Breakthrough Listen’s $\pm 4~\mathrm{Hz~s^{-1}}$ searches of nearby stars (Enriquez2017BL692), exoplanet hosts (Price2020BL1327), and the Galactic Center (Gajjar2021GC), facilitating direct comparability of sensitivity and candidate rates across diverse astrophysical contexts, now extended to GCs.
4 Results
4.1 TurboSETI Hit Statistics
Using the configuration in §3 (threshold ${\rm S/N}=10$, $\Delta\nu\simeq 7.5$ Hz, $|\dot{\nu}|\leq 4$ Hz s-1), turboSETI reported a total of $2.75\times 10^{5}$ raw hits across the five targets and both polarizations (XX: 143,890; YY: 130,875; Table 3). XX and YY totals agree within $\sim 10\%$, indicating that most interferers are not strongly polarized. Frequency histograms show pronounced concentrations within known aviation (SSR 1030–1140 MHz) and GNSS allocations (e.g., 1176/1207/1227/1268/1381 MHz), consistent with the site environment described in §4.3.
@@ 4.1 TurboSETI Hit Statistics
4.2 MBPS Filtering Outcomes and Null Result
We apply the verification framework of §3.2 to the per-beam hit lists from §3.1. In the operational order of the pipeline, the filtering proceeds by enforcing (C1)–(C5): (C1) gOTW gating per BG and pass (restricting candidate formation to illuminated intervals), then (C2) the array-wide simultaneity veto using reference beams (outside their gOTWs), followed by (C3) in-stripe ordering and timing, (C4) a single drift-rate solution with a plausible beam-response envelope, and (C5) cross-pass consistency where a position is illuminated in both passes.
@@ 4.2 MBPS Filtering Outcomes and Null Result
4.3 RFI Environment at FAST
The hit density peaks within well-known L-band allocations: aviation SSR/transponder (1030–1140 MHz) and multiple GNSS carriers (e.g., 1176, 1207, 1227, 1268, 1381 MHz). We also observe telemetry/radar activity near 1250–1300 MHz and downlink spillover toward 1.38–1.40 GHz. Representative cases (array-wide GNSS tones, out-of-sequence drifters, short bursts, and narrow instrumental lines) are rejected by the MBPS criteria in §3.2.
@@ 4.3 RFI Environment at FAST
Notes. Values are in $10^{15}$ W (total EIRP for unresolved narrowband tones integrated over a single 7.5 Hz channel). Distances $d$ from the Harris catalog (2010 edition) as cited in the manuscript. Single-crossing thresholds adopt ${\rm SEFD}=1.5$ Jy, ${\rm S/N}=10$, $N_{\rm pol}=2$, $\Delta\nu\simeq 7.5$ Hz, $t_{\rm int}\simeq 60$ s, yielding $S_{\min}\simeq 0.50$ Jy and ${\rm EIRP}_{\min}=4\pi d^{2}S_{\min}\Delta\nu$ (see §3.3). Stacked values assume non-coherent combination across up to five illuminated crossings within a BG stripe (equal weights; unmasked), i.e., $S_{\min,5}=S_{\min}/\sqrt{5}$; where geometry provides fewer crossings, sensitivity scales as $\sqrt{N_{\rm cross}}$. When both passes illuminate the same sky position, additional crossings can be combined under the same rule (see §3.2).
5 Discussion
5.1 Technosignature prevalence in metal-rich globular clusters

### v3

@@ I Introduction
The search for technosignatures—observable indicators of technology developed by extraterrestrial intelligence—has long emphasized radio transmissions, particularly ultra–narrowband features (Hz–level) that are widely treated as hallmarks of engineered emitters rather than natural astrophysical processes. This logic underpins many modern radio SETI programs, which prioritize searches for narrowband carriers across GHz bands (e.g., Siemion et al., 2013; Enriquez et al., 2017; Price et al., 2020).
@@
Despite decades of effort, no confirmed extraterrestrial signal has been identified. A central challenge is the ubiquity of terrestrial radio–frequency interference (RFI), which can closely mimic sky–localized technosignatures and inflate candidate lists by orders of magnitude before vetting (e.g., Enriquez et al., 2017; Price et al., 2020). These experiences motivate observing strategies and pipelines that discriminate RFI in situ and at scale.
@@
To that end, a MultiBeam Point-Source Scanning (MBPS) strategy has been proposed and tested with the Five-hundred-meter Aperture Spherical radio Telescope (FAST). MBPS exploits a multibeam receiver and controlled slews so that a sky-localized signal appears only in the beam presently aligned with the source, whereas broad RFI typically enters multiple beams simultaneously. In a FAST trial, MBPS logic rejected all initial narrowband candidates as RFI, illustrating the approach’s power to suppress persistent false positives (Huang et al., 2023).
@@
FAST is exceptionally well suited to such strategies. Its 19-beam L-band receiver (1.05–1.45 GHz) and low system temperature deliver outstanding sensitivity for technosignature searches (Jiang et al., 2020a; Nan et al., 2011; Li, 2016). The multibeam architecture is integral to MBPS and enables simultaneous on-source/off-source comparisons within a single pointing pattern.
@@
Beyond nearby stars, there is growing interest in extending searches to diverse environments. Globular clusters (GCs) are ancient, gravitationally bound stellar systems with $10^{5}$$10^{6}$ members; their high stellar densities could shorten communication and travel times within a putative technological network, yet crowding and typically low metallicity may inhibit planet formation and long-term orbital stability (Fischer and Valenti, 2005; Beer et al., 2004; Sigurdsson et al., 2003; Di Stefano and Ray, 2016). From an observational perspective, GC demographics and structural parameters are well compiled in the Harris catalog and updates, which inform experiment design (Harris, 1996).
@@
This work presents the first FAST SETI survey dedicated to GCs and the first to employ MBPS as the primary observing strategy. We target NGC 6171 (M107), NGC 6218 (M12), NGC 6254 (M10), NGC 6838 (M71), and IC 1276 (Pal 7). Our pilot survey demonstrates MBPS in extended, dense stellar fields and places sensitive upper limits on powerful radio transmitters within these systems. Section II summarizes observations and configuration; Section III details search methods, RFI assessment, and sensitivity; Section V discusses implications and prospects for future GC technosignature searches.
II Observations
II.1 Target Selection
@@
Rationale. Globular clusters (GCs) are ancient, dense stellar systems that offer both astrophysical motivation and practical advantages for technosignature searches. Theoretical work has argued that, despite high stellar encounter rates in cluster cores, there exist substantial “sweet-spot” regions in many GCs where habitable-zone orbits can remain dynamically stable over gigayear timescales; the short interstellar separations in such regions would also greatly facilitate interstellar communication and travel for any advanced civilization (Di Stefano and Ray, 2016). At the same time, the well-established planet–metallicity correlation suggests that stars drawn from the metal-richer tail of the GC distribution should be more likely to host giant planets (and perhaps complex planetary architectures) than their more metal-poor counterparts (Fischer and Valenti, 2005). Empirically, planets can exist in GCs: the hierarchical triple in M4 (PSR B1620-26 with a planetary-mass companion) provides a long-discussed proof-of-concept that planetary bodies can survive, or be (re)assembled, in low-metallicity cluster environments (Sigurdsson et al., 2003; Beer et al., 2004). Together, these considerations motivate a focused SETI experiment on relatively metal-rich Milky Way GCs.
@@
    Metallicity (planet formation prior). We prioritized clusters toward the metal-rich end of the Galactic GC distribution, where the prior probability for giant-planet occurrence is higher (Fischer and Valenti, 2005). In particular, M71 and Pal 7 are among the more metal-rich Milky Way GCs, while M10/M12/M107 are of intermediate metallicity. All metallicities quoted in this paper are drawn uniformly from the Harris catalog (1996; 2010 edition) to ensure internal consistency (Harris, 1996).
@@
    Proximity (EIRP sensitivity). For a flux-limited narrowband search, the minimum detectable equivalent isotropic radiated power scales as $d^{2}$. We therefore favored nearby clusters (heliocentric distances $\sim$4–6.5 kpc), improving our EIRP limits for any continuous narrowband transmitters (Harris, 1996).
@@
    Stellar population size (trial factor). Integrated absolute magnitudes ($M_{V}$) from the Harris catalog were used as proxies for total stellar population. Brighter clusters (e.g., M10, M12, M107 with $M_{V}\sim-7$) host larger numbers of stars, increasing the chance of an n=1n{=}1 technosignature source within the beam at any instant (Harris, 1996).
@@
Scientific expectations. From an SETI perspective, this sample spans a scientifically useful range of cluster environments while keeping the priors favorable. The inclusion of comparatively metal-rich disk/bulge clusters (M71, Pal 7) is directly aligned with planet-occurrence evidence (Fischer and Valenti, 2005), and the presence of the PSR B1620-26 planetary system in M4 demonstrates that planetary companions (and complex dynamical assembly channels) are realizable in GCs (Sigurdsson et al., 2003; Beer et al., 2004). Theoretical work indicates that, even in dynamically active systems, large volumes outside the densest cores can host long-lived habitable zones (Di Stefano and Ray, 2016). Observationally, all five targets are close enough that FAST’s sensitivity yields stringent EIRP limits for continuous, narrowband carriers (see §III.3), and their declinations place them in FAST’s high-gain sky. For reproducibility and uniformity, the equatorial coordinates, distances, metallicities, and $M_{V}$ values used here are taken from the Harris (2010 edition) catalog (Harris, 1996); the specific parameter values and the observation windows are listed in Table 1 and Table 2.
Table 1: Target Globular Clusters and Key Properties (from Harris, 1996)
@@
Note. — Harris (1996; 2010 edition) values. Distances are heliocentric; metallicities in the Zinn–West scale. Angular sizes are from Dreyer and Sinnott (1989)
@@
II.2 FAST Instrument, MBOTF Scanning, and Geometry
@@
and the separation between adjacent on–target beam centers ($\sim$5.74′) implies $\sim$115 s between consecutive crossing midpoints at the commanded scan rate. These deterministic spacings (in angle and in time), together with the two-pass complementary coverage, are the geometric constraints that underpin our later multibeam verification (§III.2).
For subsequent analysis across the full 19-beam footprint, the array tracks four adjacent declination stripes during the raster. We therefore refer to four beam groups (BGs), A–D, each denoting one stripe traced by a subset of beams in a pass. The central stripe (BG–C) contains the five-beam line that passes the cluster core; the adjacent stripes (BG–A/B/D) provide complementary illumination above and below BG–C. This stripe-based view is purely geometric and does not change the instrument configuration; it serves to organize timing/order expectations and coverage accounting in the cross–verification stage (§III.2).
III Data Analysis
III.1 Primary Search Configuration and Gating
@@
Spectra are analyzed at $\sim$7.5 Hz frequency resolution and $\sim$10 s integration. Linear polarizations (XX and YY) are processed independently over a drift-rate range $|{\dot{\nu}}|\leq 4$ Hz s-1, and events with ${\rm S/N}\geq 10$ are recorded. The resulting per-beam hit lists contain the BG label, pass index, gOTW identifier, boresight offset from the mask, and boundary flags marking hits within a few seconds of gOTW edges. Geometry-based validation is applied afterward (§III.2). A conservative HPBW corresponding to the widest beam in 1.05–1.45 GHz is used to ensure uniform gating across the band.
III.2 MBPS Cross-Verification Using Beam Groups
@@ III.3 Sensitivity and EIRP
III.4 Choice of Maximum Drift Rate
@@
    Astrophysical plausibility. Earth’s orbital and rotational accelerations yield drifts $\lesssim 0.2~\mathrm{Hz~s^{-1}}$ at L-Band (Siemion et al., 2013). A transmitter on an Earth-like planet would plausibly add a similar contribution, placing realistic drifts in the $\sim 0.3$$0.5~\mathrm{Hz~s^{-1}}$ range. Our ceiling exceeds this by nearly an order of magnitude, comfortably encompassing Earth-like and moderately more extreme kinematics.
@@
    Algorithmic feasibility. With $\Delta\dot{\nu}=0.125~\mathrm{Hz~s^{-1}}$, $\pm 4~\mathrm{Hz~s^{-1}}$ implies $\sim 65$ trials per scan, a computationally efficient load for multi-hour FAST datasets. Extending to the “200 nHz” normalized envelope ($\dot{\nu}_{\max}\!\sim\!280~\mathrm{Hz~s^{-1}}$ at 1.4 GHz; Sheikh et al. 2020) would inflate trials by a factor of $\sim 70$, substantially increasing runtime and false-positive candidates without clear gains for 60 s scans.
@@
    Comparative precedent. Our choice aligns with Breakthrough Listen’s $\pm 4~\mathrm{Hz~s^{-1}}$ searches of nearby stars (Enriquez et al., 2017), exoplanet hosts (Price et al., 2020), and the Galactic Center (Gajjar et al., 2021), facilitating direct comparability of sensitivity and candidate rates across diverse astrophysical contexts, now extended to GCs.
IV Results
IV.1 TurboSETI Hit Statistics
Using the configuration in §III (threshold ${\rm S/N}=10$, $\Delta\nu\simeq 7.5$ Hz, $|\dot{\nu}|\leq 4$ Hz s-1), turboSETI reported a total of $2.75\times 10^{5}$ raw hits across the five targets and both polarizations (XX: 143,890; YY: 130,875; Table 3). XX and YY totals agree within $\sim 10\%$, indicating that most interferers are not strongly polarized. Frequency histograms show pronounced concentrations within known aviation (SSR 1030–1140 MHz) and GNSS allocations (e.g., 1176/1207/1227/1268/1381 MHz), consistent with the site environment described in §IV.3.
@@ IV.1 TurboSETI Hit Statistics
IV.2 MBPS Filtering Outcomes and Null Result
We apply the verification framework of §III.2 to the per-beam hit lists from §III.1. In the operational order of the pipeline, the filtering proceeds by enforcing (C1)–(C5): (C1) gOTW gating per BG and pass (restricting candidate formation to illuminated intervals), then (C2) the array-wide simultaneity veto using reference beams (outside their gOTWs), followed by (C3) in-stripe ordering and timing, (C4) a single drift-rate solution with a plausible beam-response envelope, and (C5) cross-pass consistency where a position is illuminated in both passes.
@@ IV.2 MBPS Filtering Outcomes and Null Result
IV.3 RFI Environment at FAST
The hit density peaks within well-known L-band allocations: aviation SSR/transponder (1030–1140 MHz) and multiple GNSS carriers (e.g., 1176, 1207, 1227, 1268, 1381 MHz). We also observe telemetry/radar activity near 1250–1300 MHz and downlink spillover toward 1.38–1.40 GHz. Representative cases (array-wide GNSS tones, out-of-sequence drifters, short bursts, and narrow instrumental lines) are rejected by the MBPS criteria in §III.2.
@@ IV.3 RFI Environment at FAST
Notes. Values are in $10^{15}$ W (total EIRP for unresolved narrowband tones integrated over a single 7.5 Hz channel). Distances $d$ from the Harris catalog (2010 edition) as cited in the manuscript. Single-crossing thresholds adopt ${\rm SEFD}=1.5$ Jy, ${\rm S/N}=10$, $N_{\rm pol}=2$, $\Delta\nu\simeq 7.5$ Hz, $t_{\rm int}\simeq 60$ s, yielding $S_{\min}\simeq 0.50$ Jy and ${\rm EIRP}_{\min}=4\pi d^{2}S_{\min}\Delta\nu$ (see §III.3). Stacked values assume non-coherent combination across up to five illuminated crossings within a BG stripe (equal weights; unmasked), i.e., $S_{\min,5}=S_{\min}/\sqrt{5}$; where geometry provides fewer crossings, sensitivity scales as $\sqrt{N_{\rm cross}}$. When both passes illuminate the same sky position, additional crossings can be combined under the same rule (see §III.2).
V Discussion
V.1 Technosignature prevalence in metal-rich globular clusters